# Intermediate Relation Size Bounds for Select-Project-Join Query Plans: Asymptotically Tight Characterizations

### v1

@@ 2.2. Query plans
Observe that selection is a special case of join, in particular, it corresponds to a $1$-way join. We say that two plans $p,p^{\prime}$ over the same signature $\sigma$ are semantically equivalent when, for every structure $\mathbf{D}$ over $\sigma$, it holds that $\mathsf{out}(p,\mathbf{D})=\mathsf{out}(p^{\prime},\mathbf{D})$. Further, for a set $\Sigma$ of keys over $\sigma$, we say that two plans $p,p^{\prime}$ over $\sigma$ are $\Sigma$-semantically equivalent when, for every structure $\mathbf{D}$ over $\sigma$ that satisfies $\Sigma$, it holds that $\mathsf{out}(p,\mathbf{D})=\mathsf{out}(p^{\prime},\mathbf{D})$. A subplan of a plan is defined in the usual manner. Formally, we define a set $\mathsf{subplans}(p)$, for each plan $p$, inductively as follows:
  • When $p$ has the form $R$, we define $\mathsf{subplans}(p)=\{p\}$.
  • When $p$ has the form $\sigma_{\theta}(p^{\prime})$ or $\pi_{j_{1},\ldots,j_{n}}(p^{\prime})$, we define $\mathsf{subplans}(p)=\mathsf{subplans}(p^{\prime})\cup\{p\}.$
  • When $p$ has the form $\Join_{\theta}(p_{1},\ldots,p_{\ell})$, we define $\mathsf{subplans}(p)=(\bigcup_{i=1}^{\ell}\mathsf{subplans}(p_{i}))\cup\{p\}.$
We then say that $q$ is a subplan of $p$ when $q\in\mathsf{subplans}(p)$. Towards establishing our positive results, that is, presenting SPJ plans satisfying certain upper bounds, we rely on the notion of well-behaved plans. A plan $p$ is well-behaved if each subplan $q$ that is a join $\Join_{\theta}(p_{1},\ldots,p_{\ell})$ of plans $p_{1},\ldots,p_{\ell}$ having arities $m_{1},\ldots,m_{\ell}$, letting $s$ denote the sum $m_{1}+\cdots+m_{\ell}$, the following is satisfied: there exists an index $k\in[s]$ such that for each index $i\in[s]\setminus([m_{1}]\cup\{k\})$, there exists an index $j\in[m_{1}]\cup\{k\}$ such that $(i=j)\in\theta$. In words, each join adds at most one column of new information to the plan $p_{1}$: letting $J$ denote the set containing (1) the indices in $[m_{1}]$, which correspond to $p_{1}$, and (2) another chosen index $k$; it must be that every index $i\notin J$ is equated with an index in $J$. This property allows for the efficient computation of such joins and, thus, well-behaved plans; we will make this formal in Proposition 3.4.
2.3. Hypergraphs and tree decompositions
@@
A hypergraph is a pair $(V,E)$ where $V$ is a set, called the vertex set, and $E\subseteq 2^{V}$ is a set called the edge set. A graph is a hypergraph where each edge has size $2$. When $H$ is a hypergraph, we use $V(H)$ and $E(H)$ to denote its vertex set and edge set, respectively. An isolated vertex of a hypergraph $H$ is a vertex not occurring in any edge, that is, an element of $V(H)\setminus\bigcup_{e\in E(H)}e$. When $H$ is a hypergraph and $S\subseteq V(H)$, we use $H[S]$ to denote the induced subhypergraph of $H$ on $S$, that is, the hypergraph with vertex set $S$ and with edge set $\{e\cap S~{}|~{}e\in E(H)\}$.
When $\mathbf{A}$ is a structure over signature $\sigma$ we use $\mathbf{H}(\mathbf{A})$ to denote the hypergraph with vertex set $\mathsf{dom}(\mathbf{A})$ and edge set $\{\{\overline{a}\}~{}|~{}R\in\sigma,\overline{a}\in R^{\mathbf{A}}\}$. When $(\mathbf{A},\overline{a})$ is an open structure, we use $\mathbf{H}(\mathbf{A},\overline{a})$ to denote the hypergraph with vertex set $\mathsf{dom}(\mathbf{A})$ and with edge set $E(\mathbf{H}(\mathbf{A}))\cup\{\{\overline{a}\}\}$.
A tree decomposition of a hypergraph $H$ is a pair $(T,\chi)$, where $T$ is a tree and $\chi$ is a labeling function $V(T)\rightarrow 2^{V(H)}$, i.e., it assigns each vertex of $T$ a subset of vertices of $H$, such that
  1. (1)
    (vertex coverage) for each $v\in V(H)$, there exists $t\in V(T)$ such that $v\in\chi(t)$,
  2. (2)
    (edge coverage) for each $e\in E(H)$, there exists $t\in V(T)$ such that $e\subseteq\chi(t)$, and
  3. (3)
    (connectivity) for each $v\in V(H)$, the set $\{t\in V(T)~{}|~{}v\in\chi(t)\}$ induces a connected subtree of $T$.
Note that, if $H$ has no isolated vertices, condition $(1)$ is implied by condition $(2)$. A tree decomposition of an open structure $(\mathbf{A},\overline{a})$, is a tree decomposition of its hypergraph $\mathbf{H}(\mathbf{A},\overline{a})$.
2.4. Key constraints
A key constraint $\kappa$ over signature $\sigma$ is an expression $\textrm{key}(R)=K$, where $R/n\in\sigma$ and $K\subseteq[n]$. For an $n$-tuple $\overline{a}=(a_{1},\ldots,a_{n})$ and some $K=\{i_{1},\ldots,i_{m}\}\subseteq[n]$, for some $m\in[n]$, we write $\overline{a}[K]$ for the tuple $(a_{i_{1}},\ldots,a_{i_{m}})$. A structure $\mathbf{A}$ over signature $\sigma$ satisfies $\kappa$ if, for every two tuples $\overline{a},\overline{b}\in R^{\mathbf{A}}$, $\overline{a}[K]=\overline{b}[K]$ implies $\overline{a}=\overline{b}$. If, for some key $\kappa$ of the form $\textrm{key}(R)=K$, the set $K$ is a singleton, we say that $\kappa$ is a unary key. We say a structure satisfies a set of keys $\Sigma$, if it satisfies every key $\kappa\in\Sigma$.
The chase is a well-known tool for enforcing key dependencies (see, e.g., (Maier et al., 1979; Deutsch et al., 2006)). For an open structure $(\mathbf{A},\overline{a})$ over signature $\sigma$ and a set $\Sigma$ of keys over $\sigma$ a chase step on $(\mathbf{A},\overline{a})$ w.r.t. $\Sigma$ is a substitution of the form $x\mapsto y$, which is applied globally on $\mathbf{A}$ and $\overline{a}$, where there are two tuples $\overline{b},\overline{b^{\prime}}\in R^{\mathbf{A}}$ and a key $\textrm{key}(R)=K$ for some $R/n\in\sigma$, such that $\overline{b}[K]=\overline{b^{\prime}}[K]$ but there exists some $i\in[n]\setminus K$ such that $\overline{b}[i]=x$ and $\overline{b^{\prime}}[i]=y$. The chase of $(\mathbf{A},\overline{a})$ w.r.t. $\Sigma$, denoted $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})$, is the result of repeated applications of chase steps until a fixpoint is reached; this fixpoint is known to be unique up to isomorphism. The chase of a structure $\mathbf{A}$ over signature $\sigma$ w.r.t. $\Sigma$, denoted $\textrm{Chase}_{\Sigma}(\mathbf{A})$, is defined analogously in the obvious way.
We next define a measure for hypergraphs based on a notion, called color number, defined on conjunctive queries and introduced in (Gottlob et al., 2012). For a structure $\mathbf{A}$ over signature $\sigma$ and a set $\Sigma$ of keys over $\sigma$, a valid coloring of $\mathbf{A}$ w.r.t. $\Sigma$ is a mapping $\mathsf{Col}_{\Sigma}^{\mathbf{A}}:\mathsf{dom}(\mathbf{A})\rightarrow 2^{\{1,\ldots,k\}}$, for some $k>0$, which assigns to each element of the universe of $\mathbf{A}$ a subset of $k$ colors such that the following hold:
  1. (1)
    For each key $\textrm{key}(R)=\{i_{1},\ldots,i_{m}\}\in\Sigma$, i.e., $R/n\in\sigma$ and $\{i_{1},\ldots,i_{m}\}\subseteq[n]$, we have that for every $\overline{a}\in R^{\mathbf{A}}$ and every $i\in[n]\setminus\{i_{1},\ldots,i_{m}\}$ it holds that $\mathsf{Col}_{\Sigma}^{\mathbf{A}}(\overline{a}[i])\subseteq\bigcup_{j\in[m]}\mathsf{Col}_{\Sigma}^{\mathbf{A}}(\overline{a}[i_{j}])$.
  2. (2)
    There exists an element $x\in\mathsf{dom}(\mathbf{A})$ such that $\mathsf{Col}_{\Sigma}^{\mathbf{A}}(x)\neq\emptyset$.
For a structure $\mathbf{A}$ over signature $\sigma$, a set $S\subseteq\mathsf{dom}(\mathbf{A})$, and a set $\Sigma$ of keys over $\sigma$ the color number of $S$ w.r.t. $\mathbf{A}$ and $\Sigma$, denoted $C_{\Sigma}^{\mathbf{A}}(S)$, is defined as the maximal ratio of colors assigned to the elements in $S$ over the maximum number of colors assigned to the elements of some tuple in $\mathbf{A}$ which is attainable by any valid coloring (for any $k$) of $\mathbf{A}$ w.r.t. $\Sigma$. Formally, where $\max_{\mathsf{Col}_{\Sigma}^{\mathbf{A}}}$ ranges over all such valid colorings:
$C_{\Sigma}^{\mathbf{A}}(S)=\max_{\mathsf{Col}_{\Sigma}^{\mathbf{A}}}\frac{|\bigcup_{x\in S}\mathsf{Col}_{\Sigma}^{\mathbf{A}}(x)|}{\max_{\overline{b}\in\mathbf{A}}|\bigcup_{x\in\overline{b}}\mathsf{Col}_{\Sigma}^{\mathbf{A}}(x)|}.$
A crucial observation is that for a coloring that maximises the color number of $S$, we have that $x\notin S$ implies $\mathsf{Col}_{\Sigma}^{\mathbf{A}}(x)=\emptyset$. Note that, for a structure $\mathbf{A}$ over signature $\sigma$, the color number $C_{\Sigma}^{\mathbf{A}}(S)$, as defined here, is equal to the color number, as defined in (Gottlob et al., 2012), of the conjunctive query whose atoms are those in $\{R(v_{1},\ldots,v_{k})~{}|~{}R\in\sigma,(v_{1},\ldots,v_{k})\in R\}$, and whose output variables are $S$. We refer the reader to (Gottlob et al., 2012) for more information on the color number.
@@ 2.4. Key constraints
Let us view $C_{\Sigma}^{\mathbf{A}}$ as a mapping from the power set of $\mathsf{dom}(\mathbf{A})$ to $\mathbb{Q}^{+}$. We define the following width notions. For an open structure $(\mathbf{A},\overline{a})$ over signature $\sigma$, a set $\Sigma$ of keys over $\sigma$, and a tree decomposition $F=(T,\chi)$ of $\mathbf{H}(\mathbf{A},\overline{a})$, let $C_{\Sigma}^{\mathbf{A}}\textup{-width}(F)=\max\{C_{\Sigma}^{\mathbf{A}}(\chi(t))~{}|~{}t\in V(T)\}$. Further, let $C_{\Sigma}^{\mathbf{A}}\textup{-width}(\mathbf{A},\overline{a})=\min_{F}C_{\Sigma}^{\mathbf{A}}\textup{-width}(F)$, where $F$ ranges over all tree decompositions of $\mathbf{H}(\mathbf{A},\overline{a})$.
3. Intermediate Degree and Main Theorem Statement

### v2

@@ 2.2 Query plans
For completeness and clarity, we precisely define the types of plans to be studied. For each $m\geq 0$, we define an $m$-suitable identification as an expression $j=k$ where $j,k\in[m]$. For a signature $\sigma$, a SPJU plan over $\sigma$ is inductively defined as follows: (basic) For each $R\in\sigma$, it holds that $R$ is a plan of arity $\mathrm{ar}(R)$. (select) When $p$ is a plan of arity $m$ and $\theta$ is a set of $m$-suitable identifications, $\sigma_{\theta}(p)$ is a plan of arity $m$. (project) When $p$ is a plan of arity $m$ and $j_{1},\ldots,j_{n}$ is a sequence of numbers from $[m]$, with $n\geq 0$, it holds that $\pi_{j_{1},\ldots,j_{n}}(p)$ is a plan of arity $n$. (join) When $p_{1},\ldots,p_{\ell}$ are plans of arity $m_{1},\ldots,m_{\ell}$, respectively, with $\ell\geq 1$, and $\theta$ is a set of $(m_{1}+\cdots+m_{\ell})$-suitable identifications, $\Join_{\theta}(p_{1},\ldots,p_{\ell})$ is a plan of arity $(m_{1}+\cdots+m_{\ell})$. (union) When $p_{1},p_{2}$ are plans of arity $m$, it holds that $p_{1}\cup p_{2}$ is a plan of arity $m$. In this paper, we will sometimes simply use the term plan to refer to an SPJU plan. We define an SPJ plan to be a SPJU plan that does not make use of union. Suppose that $Q$ is a relation of arity $m$. For any set $\theta$ of $m$-suitable identifications, we define $\sigma_{\theta}(Q)=\{(d_{1},\ldots,d_{m})\in Q\penalty 10000\ |\penalty 10000\ \mbox{for each $(j=k)\in\theta$, $d_{j}=d_{k}$ holds}\}.$ For any sequence $j_{1},\ldots,j_{n}\in[m]$, we define $\pi_{j_{1},\ldots,j_{n}}(Q)=\{(d_{j_{1}},\ldots,d_{j_{n}})\penalty 10000\ |\penalty 10000\ (d_{1},\ldots,d_{m})\in Q\}.$ Let $p$ be a SPJU plan over a signature $\sigma$, let $\mathbf{D}$ be a structure over $\sigma$, and let $m$ be the arity of $p$. We define the output of $p$ over $\mathbf{D}$, denoted by $\mathsf{out}(p,\mathbf{D})$, to be the arity $m$ relation over $D$ defined as follows. Note that, as is usual in this context, we understand that there is a unique tuple of arity 0, called the empty tuple. When $p$ has the form $R$, we define $\mathsf{out}(p,\mathbf{D})=R^{\mathbf{D}}.$ When $p$ has the form $\sigma_{\theta}(p^{\prime})$, we define $\mathsf{out}(p,\mathbf{D})=\sigma_{\theta}(\mathsf{out}(p^{\prime},\mathbf{D})).$ When $p$ has the form $\pi_{j_{1},\ldots,j_{n}}(p^{\prime})$, we define $\mathsf{out}(p,\mathbf{D})=\pi_{j_{1},\ldots,j_{n}}(\mathsf{out}(p^{\prime},\mathbf{D})).$ When $p$ has the form $\Join_{\theta}(p_{1},\ldots,p_{\ell})$, we define $\mathsf{out}(p,\mathbf{D})=\sigma_{\theta}(\mathsf{out}(p_{1},\mathbf{D})\times\cdots\times\mathsf{out}(p_{\ell},\mathbf{D})).$ When $p$ has the form $p_{1}\cup p_{2}$, we define $\mathsf{out}(p,\mathbf{D})=\mathsf{out}(p_{1},\mathbf{D})\cup\mathsf{out}(p_{2},\mathbf{D}).$ Observe that selection is a special case of join, in particular, it corresponds to a $1$-way join. We say that two plans $p,p^{\prime}$ over the same signature $\sigma$ are semantically equivalent when, for every structure $\mathbf{D}$ over $\sigma$, it holds that $\mathsf{out}(p,\mathbf{D})=\mathsf{out}(p^{\prime},\mathbf{D})$. Further, for a set $\Sigma$ of keys over $\sigma$, we say that two plans $p,p^{\prime}$ over $\sigma$ are $\Sigma$-semantically equivalent when, for every structure $\mathbf{D}$ over $\sigma$ that satisfies $\Sigma$, it holds that $\mathsf{out}(p,\mathbf{D})=\mathsf{out}(p^{\prime},\mathbf{D})$. A subplan of a plan is defined in the usual manner. Formally, we define a set $\mathsf{subplans}(p)$, for each plan $p$, inductively as follows: When $p$ has the form $R$, we define $\mathsf{subplans}(p)=\{p\}$. When $p$ has the form $\sigma_{\theta}(p^{\prime})$ or $\pi_{j_{1},\ldots,j_{n}}(p^{\prime})$, we define $\mathsf{subplans}(p)=\mathsf{subplans}(p^{\prime})\cup\{p\}.$ When $p$ has the form $\Join_{\theta}(p_{1},\ldots,p_{\ell})$, we define $\mathsf{subplans}(p)=(\bigcup_{i=1}^{\ell}\mathsf{subplans}(p_{i}))\cup\{p\}.$ When $p$ has the form $p_{1}\cup p_{2}$, we define $\mathsf{subplans}(p)=\mathsf{subplans}(p_{1})\cup\mathsf{subplans}(p_{2})\cup\{p\}.$ We then say that $q$ is a subplan of $p$ when $q\in\mathsf{subplans}(p)$.
2.3 Hypergraphs and tree decompositions
@@
A hypergraph is a pair $(V,E)$ where $V$ is a set, called the vertex set, and $E\subseteq 2^{V}$ is a set called the edge set. A graph is a hypergraph where each edge has size $2$. When $H$ is a hypergraph, we use $V(H)$ and $E(H)$ to denote its vertex set and edge set, respectively. An isolated vertex of a hypergraph $H$ is a vertex not occurring in any edge, that is, an element of $V(H)\setminus\bigcup_{e\in E(H)}e$. When $H$ is a hypergraph and $S\subseteq V(H)$, we use $H[S]$ to denote the induced subhypergraph of $H$ on $S$, that is, the hypergraph with vertex set $S$ and with edge set $\{e\cap S\penalty 10000\ |\penalty 10000\ e\in E(H)\}$. When $\mathbf{A}$ is a structure over signature $\sigma$ we use $\mathbf{H}(\mathbf{A})$ to denote the hypergraph with vertex set $\mathsf{dom}(\mathbf{A})$ and edge set $\{\{\overline{a}\}\penalty 10000\ |\penalty 10000\ R\in\sigma,\overline{a}\in R^{\mathbf{A}}\}$. When $(\mathbf{A},\overline{a})$ is an open structure, we use $\mathbf{H}(\mathbf{A},\overline{a})$ to denote the hypergraph with vertex set $\mathsf{dom}(\mathbf{A})$ and with edge set $E(\mathbf{H}(\mathbf{A}))\cup\{\{\overline{a}\}\}$. A tree decomposition of a hypergraph $H$ is a pair $(T,\chi)$, where $T$ is a tree and $\chi$ is a labeling function $V(T)\rightarrow 2^{V(H)}$, i.e., it assigns each vertex of $T$ a subset of vertices of $H$, such that 1. (vertex coverage) for each $v\in V(H)$, there exists $t\in V(T)$ such that $v\in\chi(t)$, 2. (edge coverage) for each $e\in E(H)$, there exists $t\in V(T)$ such that $e\subseteq\chi(t)$, and 3. (connectivity) for each $v\in V(H)$, the set $\{t\in V(T)\penalty 10000\ |\penalty 10000\ v\in\chi(t)\}$ induces a connected subtree of $T$. Note that, if $H$ has no isolated vertices, condition $(1)$ is implied by condition $(2)$. A tree decomposition of an open structure $(\mathbf{A},\overline{a})$, is a tree decomposition of its hypergraph $\mathbf{H}(\mathbf{A},\overline{a})$.
2.4 Key constraints
A key constraint $\kappa$ over signature $\sigma$ is an expression $\textrm{key}(R)=K$, where $R/n\in\sigma$ and $K\subseteq[n]$. For an $n$-tuple $\overline{a}=(a_{1},\ldots,a_{n})$ and some $K=\{i_{1},\ldots,i_{m}\}\subseteq[n]$, for some $m\in[n]$, we write $\overline{a}[K]$ for the tuple $(a_{i_{1}},\ldots,a_{i_{m}})$. A structure $\mathbf{A}$ over signature $\sigma$ satisfies $\kappa$ if, for every two tuples $\overline{a},\overline{b}\in R^{\mathbf{A}}$, $\overline{a}[K]=\overline{b}[K]$ implies $\overline{a}=\overline{b}$. If, for some key $\kappa$ of the form $\textrm{key}(R)=K$, the set $K$ is a singleton, we say that $\kappa$ is a unary key. We say a structure satisfies a set of keys $\Sigma$, if it satisfies every key $\kappa\in\Sigma$. The chase is a well-known tool for enforcing key dependencies (see, e.g., [16, 8]). For an open structure $(\mathbf{A},\overline{a})$ over signature $\sigma$ and a set $\Sigma$ of keys over $\sigma$ a chase step on $(\mathbf{A},\overline{a})$ w.r.t. $\Sigma$ is a substitution of the form $x\mapsto y$, which is applied globally on $\mathbf{A}$ and $\overline{a}$, where there are two tuples $\overline{b},\overline{b^{\prime}}\in R^{\mathbf{A}}$ and a key $\textrm{key}(R)=K$ for some $R/n\in\sigma$, such that $\overline{b}[K]=\overline{b^{\prime}}[K]$ but there exists some $i\in[n]\setminus K$ such that $\overline{b}[i]=x$ and $\overline{b^{\prime}}[i]=y$. The chase of $(\mathbf{A},\overline{a})$ w.r.t. $\Sigma$, denoted $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})$, is the result of repeated applications of chase steps until a fixpoint is reached; this fixpoint is known to be unique up to isomorphism. The chase of a structure $\mathbf{A}$ over signature $\sigma$ w.r.t. $\Sigma$, denoted $\textrm{Chase}_{\Sigma}(\mathbf{A})$, is defined analogously in the obvious way. We next define a measure for hypergraphs based on a notion, called color number, defined on conjunctive queries and introduced in [10]. For a structure $\mathbf{A}$ over signature $\sigma$ and a set $\Sigma$ of keys over $\sigma$, a valid coloring of $\mathbf{A}$ w.r.t. $\Sigma$ is a mapping $\mathsf{Col}_{\Sigma}^{\mathbf{A}}:\mathsf{dom}(\mathbf{A})\rightarrow 2^{\{1,\ldots,k\}}$, for some $k>0$, which assigns to each element of the universe of $\mathbf{A}$ a subset of $k$ colors such that the following hold: 1. For each key $\textrm{key}(R)=\{i_{1},\ldots,i_{m}\}\in\Sigma$, i.e., $R/n\in\sigma$ and $\{i_{1},\ldots,i_{m}\}\subseteq[n]$, we have that for every $\overline{a}\in R^{\mathbf{A}}$ and every $i\in[n]\setminus\{i_{1},\ldots,i_{m}\}$ it holds that $\mathsf{Col}_{\Sigma}^{\mathbf{A}}(\overline{a}[i])\subseteq\bigcup_{j\in[m]}\mathsf{Col}_{\Sigma}^{\mathbf{A}}(\overline{a}[i_{j}])$. 2. There exists an element $x\in\mathsf{dom}(\mathbf{A})$ such that $\mathsf{Col}_{\Sigma}^{\mathbf{A}}(x)\neq\emptyset$. For a structure $\mathbf{A}$ over signature $\sigma$, a set $S\subseteq\mathsf{dom}(\mathbf{A})$, and a set $\Sigma$ of keys over $\sigma$ the color number of $S$ w.r.t. $\mathbf{A}$ and $\Sigma$, denoted $C_{\Sigma}^{\mathbf{A}}(S)$, is defined as the maximal ratio of colors assigned to the elements in $S$ over the maximum number of colors assigned to the elements of some tuple in $\mathbf{A}$ which is attainable by any valid coloring (for any $k$) of $\mathbf{A}$ w.r.t. $\Sigma$. Formally, where $\max_{\mathsf{Col}_{\Sigma}^{\mathbf{A}}}$ ranges over all such valid colorings: $C_{\Sigma}^{\mathbf{A}}(S)=\max_{\mathsf{Col}_{\Sigma}^{\mathbf{A}}}\frac{|\bigcup_{x\in S}\mathsf{Col}_{\Sigma}^{\mathbf{A}}(x)|}{\max_{\overline{b}\in\mathbf{A}}|\bigcup_{x\in\overline{b}}\mathsf{Col}_{\Sigma}^{\mathbf{A}}(x)|}.$ A crucial observation is that for a coloring that maximises the color number of $S$, we have that $x\notin S$ implies $\mathsf{Col}_{\Sigma}^{\mathbf{A}}(x)=\emptyset$. Note that, for a structure $\mathbf{A}$ over signature $\sigma$, the color number $C_{\Sigma}^{\mathbf{A}}(S)$, as defined here, is equal to the color number, as defined in [10], of the conjunctive query whose atoms are those in $\{R(v_{1},\ldots,v_{k})\penalty 10000\ |\penalty 10000\ R\in\sigma,(v_{1},\ldots,v_{k})\in R\}$, and whose output variables are $S$. We refer the reader to [10] for more information on the color number, but note the following fact here.
Proposition 2.3.
(refer to [10, Section 3]) Let $\mathbf{B}$ be a structure, and let $S$ be a subset of $\mathsf{dom}(\mathbf{B})$. The value $C^{\mathbf{B}}_{\emptyset}(S)$ is the fractional edge cover number of the induced hypergraph of $\mathbf{H}(\mathbf{B})$ on $S$.
@@ 2.4 Key constraints
A fractional edge cover of a hypergraph $G=(V,F)$ is a mapping $x:F\to[0,\infty)$ such that for each $v\in V$, it holds that $\sum_{e\in E,v\in e}x(e)\geq 1$; the sum $\sum_{e\in E}x(e)$ is the weight of $x$. The fractional edge cover number of $G$ is the minimum weight over all fractional edge covers of $G$; it follows from the theory of linear programming that a minimum exists [12]. Let us view $C_{\Sigma}^{\mathbf{A}}$ as a mapping from the power set of $\mathsf{dom}(\mathbf{A})$ to $\mathbb{Q}^{+}$. We define the following width notions. For an open structure $(\mathbf{A},\overline{a})$ over signature $\sigma$, a set $\Sigma$ of keys over $\sigma$, and a tree decomposition $F=(T,\chi)$ of $\mathbf{H}(\mathbf{A},\overline{a})$, let $C_{\Sigma}^{\mathbf{A}}\textup{-width}(F)=\max\{C_{\Sigma}^{\mathbf{A}}(\chi(t))\penalty 10000\ |\penalty 10000\ t\in V(T)\}$. Further, let $C_{\Sigma}^{\mathbf{A}}\textup{-width}(\mathbf{A},\overline{a})=\min_{F}C_{\Sigma}^{\mathbf{A}}\textup{-width}(F)$, where $F$ ranges over all tree decompositions of $\mathbf{H}(\mathbf{A},\overline{a})$.
3 Degree Notions and Main Theorem Statement
3.1 Degree notions